OpenStudy (anonymous):

The area of a rectangle is 27m^2, and the length of the rectangle is 3 m less than twice the width. Find the dimension of the rectangle.

OpenStudy (nikato):

let w=width and we know the length of the rectangle is 3 m less than twice the width so that means length=2w-3

OpenStudy (nikato):

area of a rectangle is 27m^2 means length x width= 27 so w(2w-3)=27 and can you solve for w from here?

OpenStudy (anonymous):

are we distrubuting?

OpenStudy (nikato):

correct. distribute the left side first

OpenStudy (anonymous):

is it 2w^2-3w=27?

OpenStudy (nikato):

yes. now subtract 27 on both sides and you'll have a quadratic

OpenStudy (anonymous):

2w^2-3w=27=0?

OpenStudy (anonymous):

oops, i mean 2w^2-27= 3w-27... i guess something like that, this is all new to me.

OpenStudy (nikato):

 you'll get this. and do you know how to factor this?